• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 00:47:54 on March31,2005

    Entry number 142841

    keyword=Q1 tripped and reset

    The Left Q1 tripped at 00:15 due to a drop in the liquid levels.  We went to 
    controlled access to reset the magnet, and it is ramping up now.  We should be 
    ready to take data in a few minutes.